The three finalist in a contest are brought to the center of a large flat field. calculate the following displacements:
72.4 m, 32^o east of south
57.3 m, 36^o south of west
17.8 m straight south
What is the magnitude and direction of the resultant?
1
Expert's answer
2020-08-10T19:53:55-0400

"A(x,y)=(72.4\sin32°,72.4\cos32°)=(38.4m,-61.4m)"


"B(x,y)=(57.3\cos36°,57.3\sin36°)=(-46.4m,-33.7m)"


"C(x,y)=(0m,-17.8m)"


"\sum{x}=38.4-46.4=-8m"


"\sum{y}=-61.4-33.7-17.8=-112.9m"


"Displacement=\sqrt{8^2+112.9^2}=113.2(m)"


"Direction=\tan^{-1}{\frac{112.9}{8}}=85.9°" south of west.